A simple random sample of 60 items resulted in a sample mean of 65. The population standard deviation is 17.

a. Compute the 95% confidence interval for the population mean (to 1 decimal).

(  ,  )

b. Assume that the same sample mean was obtained from a sample of 120 items. Provide a 95% confidence interval for the population mean (to 2 decimals).

(  ,  )

c. What is the effect of a larger sample size on the margin of error?
Select

It increases

It decreases

It stays the same

It cannot be determined from the given data
Answer #1

Solution: Given that x = 65 8 = 17 n = 60 C = 95% = 0,95 At 95% confidence intelval the z c = 95% = 0.95 a = -c=1-c =1-0.95 -b) Given that x = 65 8 = 17 n = 120 C - 95 % = 0.95 At 95 % confidence interval the z c = 95 % = 0.95 x - 1-c -1-0.95 -0.05 e